OpenStudy (anonymous):

There is an inverse relationship between x and y. If y is 20 when x is 2, what is x when y is 8. 5 8 40 80

OpenStudy (anonymous):

@Jaynator495

jaynator495 (jaynator495):

d

OpenStudy (anonymous):

A line has a vertical change per unit of horizontal change of -5/4 and passes through point (2,2) which is another point on the line? (-3,6) (-10,8) (8,-10) (6,-3) @Jaynator495

jaynator495 (jaynator495):

im no good at that sorry :(

OpenStudy (anonymous):

Could you tag anyone you know?

OpenStudy (anonymous):

@aum

OpenStudy (aum):

If x and y have an inverse relationship it implies the product of x and y is a constant. That is, \(x_1*y_1 = x_2 * y_2\). Thus, \(2 * 20 = x_2 * 8\). Find \(x_2\).

OpenStudy (anonymous):

C?

OpenStudy (aum):

Can you show the steps?

OpenStudy (anonymous):

I have no idea how to do this

OpenStudy (aum):

Simplify the left hand side. What do you get?

OpenStudy (anonymous):

I don't get any of it.

OpenStudy (aum):

Sorry, I can't just give out answers.

OpenStudy (anonymous):

Show me how to work t out. I'm only n 5th grade

OpenStudy (aum):

The left hand side is 2 multiplied by 20. You should be able to multiply two numbers.

OpenStudy (anonymous):

40

OpenStudy (aum):

Yes. \(2 * 20 = x_2 * 8 \\ 40 = 8 * x_2 \\ \) Divide both sides by 8: \(5 = x_2\) So the answer is 5.

OpenStudy (anonymous):

Wait.. i need help one the second question. Not the first

OpenStudy (aum):

The answer posted for the first question by the previous person is wrong.

OpenStudy (anonymous):

Yeah I know. I needed help on the second question

OpenStudy (aum):

"A line has a vertical change per unit of horizontal change of -5/4" is just another way of saying the slope of the line m = -5/4 The general equation of a line in slope-intercept form is: y = mx + b Put m = -5/4 y = -5/4 * x + b ----- (1) This line passes through the point (2,2). Put x = 2 and y = 2 in equation (1) and solve for b first.

OpenStudy (anonymous):

Wooo.. that make no sense to me

OpenStudy (aum):

Tell me the part you want to be explained. If you are trying to solve this problem by now you should be familiar with slope, y-intercept, equation of a line in slope-intercept form, etc.

OpenStudy (aum):

If you are given the x and y coordinates of two points, do you know how to find the slope of the line passing through those two points?

OpenStudy (aum):

Then you have to go through your notes or book and catchup on basic concepts such as: slope of a line, y-intercept, equation of a line, etc. before you can understand or solve this problem.